Kailiytnrae:

If a ship's path is mapped on a coordinate grid, it follows a straight-line path of slope 3 and passes through point (2, 5). Part A: Write the equation of the ship’s path in slope-intercept form. Part B: A second ship follows a straight line, with the equation x + 3y − 6 = 0. Are these two ships sailing perpendicular to each other? Justify your answer. help please

MxxnLight:

@dontsaymyname e.e

XxAngelGirlxX:

A) Equation of the ship's path is y - 3 = 5(x - 2) y - 3 = 5x - 10 add 3 to both sides y = 5x - 7 B) (6, 25) ??? Put 6 in the equation and see id it produces 25...so we have.... y = 5(6) - 7 = 30 - 7 = 22 = point (6, 22) So...it does not pass through the point (6,25) C) x + 5y - 15 = 0 Rearrange as 5y = -x + 15 divide both sides by 5 y = (-1/5)x + 3 Yes....the ships are sailing on perpendicular pahs because the slope in the equation of the second ship's path is the negative reciprocal of the slope in the equation of the first ship's path.

Kailiytnrae:

@XxAngelGirlxX thank you